# Trapped-ion quantum simulations for condensed-phase chemical dynamics: seeking a quantum advantage

### v3

@@ Abstract
Simulating the quantum dynamics of molecules in the condensed phase represents a longstanding challenge in chemistry. Trapped-ion quantum systems may serve as a platform for the analog-quantum simulation of chemical dynamics that is beyond the reach of current classical-digital simulation. To identify a “quantum advantage” for these simulations, performance analysis of both analog-quantum simulation on noisy hardware and classical-digital algorithms is needed. In this Review, we make a comparison between a noisy analog trapped-ion simulator and a few choice classical-digital methods on simulating the dynamics of a model molecular Hamiltonian with linear vibronic coupling. We describe several simple Hamiltonians that are commonly used to model molecular systems, which can be simulated with existing or emerging trapped-ion hardware. These Hamiltonians may serve as stepping stones toward the use of trapped-ion simulators for systems beyond the reach of classical-digital methods. Finally, we identify dynamical regimes where classical-digital simulations seem to have the weakest performance compared to analog-quantum simulations. These regimes may provide the lowest hanging fruit to exploit potential quantum advantages.
@@ Introduction
The table below summarizes the four different approaches to solving a computation/simulation problem. For analog-quantum methods, qubit (qudit) is the unit component of a quantum system, each consisting of two ($d$) states. This Review focuses on comparing the analog-quantum and digital-classical methods for simulating quantum molecular dynamics. We expect that analog-quantum simulators will be actively used for solving various practical problems [6], at least until fault-tolerant digital-quantum computers emerge, similarly to how analog-classical computers had been used for, for example, solving differential equations [46, 47], before fault-tolerant digital-classical computers were developed.[Uncaptioned image] For the problem of simulating quantum dynamics, we are interested in finding expectation values of an observable $\hat{O}$ up to time $\tau$ and accuracy $\epsilon$, or estimating quantities related to molecular behavior that can be derived from these expectation values. We assume that the parameters that define the initial state and the time evolution are given. For example, given the initial state $\rho_{0}$ and the Hamiltonian $\hat{H}$, the problem is to find $O(t):=\text{Tr}\{\hat{O}e^{-i\hat{H}t}\rho_{0}e^{i\hat{H}t}\}$ for $t=t_{1},t_{2},...,\tau$ up to accuracy $\epsilon$, or to extract quantities of interest (rate constants, branching ratios, etc.) from these $O(t)$ values. A more general form of evolution, such as the Lindblad master equation [48], may be considered for dissipative dynamics. A model for quantum dynamics will always be an approximation of the dynamics of an actual molecular system. Nonetheless, it is valuable to solve approximate models accurately in order to improve the understanding of the quantum molecular phenomena. For example, solving approximate models leads to identifying the dependence of important chemical quantities, such as the electronic-state population at equilibrium, to the parameters and structure of the Hamiltonian. Furthermore, improved understanding may lead to better design of the experiments on an actual molecular system. A plethora of classical-digital methods exist to simulate quantum dynamics; the methods can be classified into two categories. First, a classical-digital method is numerically exact if it describes the time evolution of a given Hamiltonian up to a controllable error, which can be made arbitrarily small if infinite computing resources are used. Next, a classical-digital method is approximate if it relies on a physical or mathematical simplification of the Hamiltonian dynamics, so the method does not converge to the exact solution even given infinite computational resources. For quantum molecular dynamics, typical approximations include the use of classical-dynamical heuristics and the formal expansions of the quantum-mechanical path integral around the classical trajectories of the system (trajectories that satisfy the Lagrange equations of motion). These approximations significantly lessen the requirements for computational resources, such as run time and memory, compared to those of numerically exact methods. See “The current champion: …” subsection for examples of classical-digital methods in both categories. The term “quantum advantage” is used broadly to refer to the capability of quantum computers/simulators to solve a problem that cannot be solved using conventional computers. For the challenge of simulating quantum dynamics, we offer a specific working definition. A quantum dynamics problem of finding $O(t)$ up to time $\tau$ and accuracy $\epsilon$ is classically intractable if (i) all known numerically exact classical-digital methods require prohibitively long run time or memory given reasonably large computing resources (e.g., computing cluster of a research facility) and (ii) all known approximate classical-digital methods fail to predict $O(t)$ within accuracy $\epsilon$ at some $t\in[0,\tau]$. If a quantum (analog or digital) simulator can solve such a classically intractable problem within run time that is short enough to be practical, then a quantum advantage is achieved. This definition can be straightforwardly extended to dissipative dynamics as well. For the problem of extracting quantities related to molecular behavior, many valuable predictions may be made even with errors in the estimated values of $O(t)$, as long as general trends are captured accurately. For example, electron transfer rates are estimated by the decay of donor-state population over time, which can be robust to significant errors in the state populations when the errors are unbiased and transient. Also, approximate classical-digital methods (such as semiclassical approximations) may provide acceptable estimates even if some details of the dynamics are not captured.
@@
A typical trapped-ion system consists of a linear chain of ions confined in free space. Radio frequency (rf) electromagnetic fields are supplied by the electrodes of the trap, creating a pseudo-potential that constrains the atomic positions. After the ions are sufficiently cooled using lasers, the ions’ interactions with the confinement potential and the Coulomb repulsion among the ions reach equilibrium, producing a linear chain of ions. A typical ion has many accessible electronic states, and two stable states can be exploited to establish an atom-based qubit with excellent coherence properties. Initialization of the qubit to a desired state, and measurement of the qubit state, may be performed through optical excitation with lasers. Nuclear (bosonic) degrees of freedom exist in the ion chain through its collective vibrating motion. The ion chain’s motion can be described in the language of normal modes, where each mode is approximated by a quantum harmonic oscillator. An $n$-ion chain has $n$ motional modes in each of the ion chain’s axial direction and two other perpendicular (radial) directions, total $3n$ modes. A subset of these modes can interact with the qubits to mediate the entanglement between qubits or to provide the bosonic ingredient for analog simulations. Since the normal modes are spread through the ion chain, each mode can interact with all qubits in the chain (except for qubits located at nodes of the normal mode), allowing flexibility in the structure of the Hamiltonians that can be simulated. Panel (c) of the figure below visualizes the $n=4$ radial modes of a four-ion chain. The modes are shown in the descending order of mode frequency from top to bottom, and the relative height (direction) of each arrow represents the magnitude (sign) of the eigenvector component of each normal mode. We denote the mode with equal eigenvector components of all ions [top in (c)] as the center-of-mass (CM) mode and all other modes as non-CM modes. To manipulate the qubit states and motional states of the ion chain, electromagnetic fields are applied using lasers or magnetic-field gradients. Panels (a) and (b) of the figure below show the example of using lasers to couple the leftmost ion’s qubit states to one of the motional modes [third from top in (c)]. The two-photon Raman transition between the qubit states via the excited atomic state $|e\rangle$ is induced using the global (orange) and individually addressed (red and blue) laser beams. The frequency, phase, and amplitude of the laser fields impinging on each ion can be controlled individually, allowing flexibility in defining the Hamiltonian parameters that will be simulated on the ion chain. [Uncaptioned image] The architecture of trapped-ion systems considered in this Review is a single chain of many ions, which scales up to 20-30 ions for state-of-art systems [49, 50]. Another well-known architecture is the quantum charge-coupled device (QCCD) architecture [51], which utilizes operations such as splitting an ion from a chain, shuttling an ion, and merging an ion to another chain. While the QCCD architecture has advantages such as high-fidelity qubit operations, the motional states are not preserved during the splitting and merging, and thus cannot provide the bosonic ingredient for analog simulations.
@@
Assuming that the electronic and vibrational energy states can be obtained, numerically exact methods (see Box 1 for a definition) exist for classical-digital computers to compute quantum molecular dynamics in some regimes. Examples include the multi-configuration time-dependent Hartree (MCTDH) method [55, 56, 57], the Ab Initio Multiple Spawning (AIMS) method [58, 59, 60], the hierarchical equations of motion (HEOM) approach [61, 62, 63], the quasi-adiabatic propagator path integral (QUAPI) method [64, 65, 66, 67, 68, 69, 70], the time-evolving matrix products operators (TEMPO) method [71, 72, 73], the dissipation-assisted matrix product factorization (DAMPF) method [74, 75], and the time-dependent density-matrix renormalization group (tDMRG) method [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. The tDMRG method, for example, can efficiently simulate a composite system of electronic states and bath modes when the composite system may be mapped to a simple configuration, such as one-dimensional chain [78, 79] or star geometry [84] (see Fig. 1d), through an appropriate basis transformation. For the case of one-dimensional chain, large systems (e.g., $N\sim 10^{2}$) with moderate interaction strengths (e.g., $|\kappa_{i,j,k}|\lesssim|\Delta_{i,j}|,\>|E_{i}-E_{j}|,\>\nu_{k}$) can be simulated [79]. However, the tDMRG method is either inaccurate or computationally prohibitive when a transformation to a simple configuration cannot be found, or when the interactions are strong (e.g., $|\kappa_{i,j,k}|\sim$ 10 to $10^{2}$ times $|\Delta_{i,j}|,\>|E_{i}-E_{j}|,\>\nu_{k}$), which is a regime of interest [88, 89, 90]. Other numerically exact methods also become computationally prohibitive, in terms of computational time or memory demands, when the entanglement among the components becomes significant [91, 72, 66].
@@
This Trotterization-based analog approach to quantum simulation, also suggested for simulating processes in quantum field theory [136], contrasts with the traditional approach in analog trapped-ion simulations, where interactions for each and every term in the simulated Hamiltonian is simultaneously switched on [137, 36]. While the traditional approach enabled various quantum-simulation experiments, even with hardware that does not fully support individual control of qubits (see Ref. 138 for an example of simulating quantum molecular dynamics), the parameters of the simulated Hamiltonian should meet specific constraints [^1] For example, when a quantum spin system is simulated using the traditional approach where all ions are simultaneously illuminated by a monochromatic beam, the coupling strength $\Delta_{i,j}$ between spins $i$ and $j$, mapped to ions $i$ and $j$ indexed from the left to right of the ion chain, should satisfy $\Delta_{i,j}\propto|i-j|^{-\alpha}$ ($30<\alpha<$ 3[137, 138, 36]. that do not exist for the Trotterization-based analog approach.
@@
Using single-qubit rotation (2), spin-dependent force (3), and MS interaction (4), each term in (1) can be simulated. Specifically, the terms $E_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}$ are mapped to (2), the terms $\kappa_{i,i,k}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ are mapped to (3), and the terms $\Delta_{i,j}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{j}$ ($i\neq j$) are mapped to (4). The terms $\kappa_{i,j,k}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{j}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ ($i\neq j$) are also mapped to a sequence of operations (2)-(4), as described in Ref. 149. The evolution with respect to each term in (1) by the Trotterization time step $\tau/S$ is simulated by choosing the field parameters and time duration of each mapped trapped-ion operation appropriately. By repeating the evolution for all terms and Trotterization steps, the evolution of the entire system described by the LVCM Hamiltonian is simulated.
@@
tDMRG simulations are performed by putting (5) into a one-dimensional topology, where the electronic sites and the $N$ bath modes form a linear chain, using the unitary transformation described in Refs. 79, 84. This strategy can be used to describe the dynamics efficiently for a donor-acceptor system coupled to a bath with hundreds of modes that have a moderately large reorganization energy ($\lambda\lesssim 10\Delta$). However, the computational cost grows as $\kappa\propto\sqrt{\lambda}$ increases, and as the entanglement among the components grows. In contrast, the Ehrenfest simulations can be performed at significantly lower computation cost than the tDMRG simulations, at the expense of accuracy in describing the dynamics only up to a short time compared to $\hslash/\Delta$, as the semiclassical mean-field approximations often break down at longer times.
Figure 1: Comparison of classical (tDMRG and Ehrenfest) and quantum (trapped ion) methods for simulating the model Hamiltonian (5). $\Delta=0.08679$ eV, $\nu_{k}=[0.08679+0.01240(k-1)/(N-1)$] eV for $k=1,..N$, where $N$ is the number of bath modes. For the trapped-ion method, we assume using $(\lceil N/2\rceil+1)$-ion chain to generate $N$ radial non-CM motional modes. (a) Evolution of the donor population, for the indicated reorganization energy values ($\lambda$) and $N=2$, simulated using various methods. For the curves labeled “Trapped ion”, we numerically simulate the evolution of a trapped-ion simulator system. The ideal case assumes no decoherence, and the noisy case assumes state-of-art noise parameters in Table 1. For both cases, shot noise (population uncertainty due to finite number of binary measurements) is not shown. In all panels, the blue and red solid curves overlap. (b) Computational run time for the tDMRG method, with various values of $\lambda/\Delta$ and $N$ [labels shared with (c)]. The additive error for each population value is less than $10^{-3}$ ($10^{-2}$) for the solid (dashed) curves. See Ref. 84 for details of the method (the interaction-picture chain-geometry algorithm). The run time scales nearly exponentially with $\sqrt{\lambda/\Delta}$, where the rate of increase is smaller for the dashed curves. Computational run time for the Ehrenfest method is negligible compared to the tDMRG method. (c) Estimated experimental time for the trapped-ion simulation, with various values of $\lambda/\Delta$ and $N$. We assume 40 equally-spaced time steps and 100 (2500) simulation runs per time step for the left (right) vertical axis, where each run produces a binary measurement result. Dotted line shows the total time spent for cooling, state preparation, and measurement, which is estimated to be 4.25 ms [158] times the total number of simulation runs. The estimated experimental time roughly scales linearly with $\sqrt{\lambda/\Delta}$. (d) A schematic representation of the star-geometry model simulated here, where $N=5$ [see (5)]. The grey circle, denoted ‘S’, represents the system with two electronic energy levels, and each parabola represents a bath mode.
@@
Even when a pulse is resonant with the sideband frequency of a mode ($\mu=\omega_{k}$), unwanted off-resonant excitations of other modes, also known as cross-mode coupling, may occur because of the finite spacings between the mode frequencies. Naïvely, this leads to a constraint that $\tilde{\Omega}_{p,k}$ should be much smaller than the mode-frequency spacing, which imposes a lower bound on the time duration of the corresponding operations. However, cross-mode coupling can be actively removed by modulating the field parameters during the operation. Here, we consider using a frequency($\mu$)-modulated pulse [160, 161] for each spin-dependent-force operation, such that cross-mode coupling is removed using operations with as short a time duration as possible. The required time duration increases as the number of ions increases and mode-frequency spacing decreases. For example, for the case of $\lambda=30\Delta$ and $S=600$, the length of each pulse that simulates a single Trotterization step averaged over all modes is 15.7 $\mu$s, 17.4 $\mu$s, and 19.0 $\mu$s for a 2-ion ($N=2$), 3-ion ($N=3,4$), and 4-ion ($N=5$) chain, respectively. The relevant system parameters for a state-of-art trapped-ion setup [162] are shown in Table 1.
In state-of-art trapped-ion systems, the leading sources of errors are motional-mode dephasing, motional-mode heating, and laser dephasing [162, 163, 164]. To simulate the evolution of a trapped-ion system in the presence of noise, we use Qutip [165] to solve the Lindblad master equation [48], following the method in Ref. 162 which provides reliable predictions of the system performance [162, 163]. We use the best recorded noise parameters from a state-of-art trapped-ion system [162, 163] as shown in Table 1, which are far from the fundamental limits and can, in principle, be improved.
@@
Table 1: System and noise parameters for the state-of-art trapped-ion system in Refs. 162, 163 that are used in Fig. 1. Motional modes in two radial directions of ion chain consisting of two to four ions are used. Motional-mode coherence time is taken from Ref. 162 and motional-mode heating rate and laser coherence time are taken from Ref. 163. The heating rate of CM motional modes is of the order of 100 quanta/s, so we assume that CM modes are not used in our trapped-ion simulations.
@@
Color gradients in the figure illustrate the qualitative performance (accuracy and computational cost) of three classes of quantum dynamics simulations of LVCMs: numerically exact classical-digital methods, approximate classical-digital methods, and analog trapped-ion simulations on a noisy device. The qualitative performance of these methods is illustrated over several regimes of coupling strength between the electronic states and bath modes, quantified by the ratio of the reorganization energy $\lambda$ to the coupling strength between electronic states $\Delta$. If $\lambda\ll\Delta$, the numerically exact methods are generally efficient and accurate, since the entanglement among the electronic states and bath modes is weak. In this regime, the approximate methods are also accurate, at least for a few dynamical cycles of the system. This is because the influence of the bath on the electronic states is weak, and errors associated with neglecting non-classical trajectories of the bath modes have little influence on the evolution of the electronic states. In the regime where $\lambda\gg\Delta$, the numerically exact methods converge slowly and generally perform poorly because of strong entanglement among the electronic states and bath modes. However, the approximate methods generally perform well, providing reliable predictions for the donor-state population decay rate, for example. This acceptable performance arises when the coupling is strong and the bath modes are highly-excited harmonic oscillators that often behave almost classically. In the intermediate regime where $\lambda\gtrsim\Delta$, the coupling can be sufficiently strong that large entanglement renders the numerically exact methods computationally intractable, yet weak enough that the bath modes are not sufficiently highly excited to behave classically. As such, this regime appears to be a compelling target for testing whether an analog trapped-ion simulation may outperform classical-digital algorithms for simulating quantum molecular dynamics economically, at least in the space of LVCMs. The upper bound of $\lambda/\Delta$ that may allow a trapped-ion simulation to produce sufficiently accurate predictions is determined by the magnitude of the noise relative to the strength of native operations of the trapped-ion system. [Uncaptioned image]

### v5

# Seeking a quantum advantage with trapped-ion quantum simulations of condensed-phase chemical dynamics

@@ Abstract
Simulating the quantum dynamics of molecules in the condensed phase represents a longstanding challenge in chemistry. Trapped-ion quantum systems may serve as a platform for the analog-quantum simulation of chemical dynamics that is beyond the reach of current classical-digital simulation. To identify a ‘quantum advantage’ for these simulations, performance analysis of both analog-quantum simulation on noisy hardware and classical-digital algorithms is needed. In this Review, we make a comparison between a noisy analog trapped-ion simulator and a few choice classical-digital methods on simulating the dynamics of a model molecular Hamiltonian with linear vibronic coupling. We describe several simple Hamiltonians that are commonly used to model molecular systems, which can be simulated with existing or emerging trapped-ion hardware. These Hamiltonians may serve as stepping stones toward the use of trapped-ion simulators for systems beyond the reach of classical-digital methods. Finally, we identify dynamical regimes where classical-digital simulations seem to have the weakest performance compared to analog-quantum simulations. These regimes may provide the lowest hanging fruit to exploit potential quantum advantages.
@@ Introduction
The table below summarizes the four different approaches to solving a computation/simulation problem. In the description of analog-quantum methods, qubit (qudit) is the unit component of a quantum system, each consisting of two ($d$) states. This Review focuses on comparing the analog-quantum and digital-classical methods for simulating quantum molecular dynamics. We expect that analog-quantum simulators will be actively used for solving various problems in chemistry, physics, and material science [6], at least until fault-tolerant digital-quantum computers emerge, similarly to how analog-classical computers had been used for, for example, solving differential equations [46, 47], before fault-tolerant digital-classical computers were developed.
For the problem of simulating quantum dynamics, we are interested in finding expectation values of an observable $\hat{O}$ up to time $\tau$ and accuracy $\epsilon$, or estimating quantities related to molecular behavior that can be derived from these expectation values. We assume that the parameters that define the initial state and the time evolution are given. For example, given the initial state $\rho_{0}$ and the Hamiltonian $\hat{H}$, the problem is to find $O(t):=\text{Tr}\{\hat{O}e^{-i\hat{H}t}\rho_{0}e^{i\hat{H}t}\}$ for $t=t_{1},t_{2},...,\tau$ up to accuracy $\epsilon$, or to extract quantities of interest (rate constants, branching ratios, etc.) from these $O(t)$ values. A more general form of evolution, such as the Lindblad master equation [48], may be considered for dissipative dynamics.
A model for quantum dynamics will always be an approximation of the dynamics of an actual molecular system. Nonetheless, it is valuable to solve approximate models accurately in order to improve the understanding of the quantum molecular phenomena. For example, solving approximate models leads to identifying the dependence of important chemical quantities, such as the electronic-state population at equilibrium, to the parameters and structure of the Hamiltonian. Furthermore, improved understanding may lead to better design of the experiments on an actual molecular system.
A plethora of classical-digital methods exist to simulate quantum dynamics; the methods can be classified into two categories. First, a classical-digital method is numerically exact if it describes the time evolution of a given Hamiltonian up to a controllable error, which can be made arbitrarily small if infinite computing resources are used. Next, a classical-digital method is approximate if it relies on a physical or mathematical simplification of the Hamiltonian dynamics, so the method does not converge to the exact solution even given infinite computational resources. For quantum molecular dynamics, typical approximations include the use of classical-dynamical heuristics and the formal expansions of the quantum-mechanical path integral around the classical trajectories of the system. These approximations significantly lessen the requirements for computational resources, such as run time and memory, compared to those of numerically exact methods. See “The current champion: …” subsection for examples of classical-digital methods in both categories.
The term “quantum advantage” is used broadly to refer to the capability of quantum computers/simulators to solve a problem that cannot be solved using conventional computers. For the challenge of simulating quantum dynamics, we offer a specific working definition. A quantum dynamics problem of finding $O(t)$ up to time $\tau$ and accuracy $\epsilon$ is classically intractable if (i) all known numerically exact classical-digital methods require prohibitively long run time or memory given reasonably large computing resources (e.g., computing cluster of a research facility) and (ii) all known approximate classical-digital methods fail to predict $O(t)$ within accuracy $\epsilon$ at some $t\in[0,\tau]$. If a quantum (analog or digital) simulator can solve such a classically intractable problem within run time that is short enough to be practical, then a quantum advantage is achieved. This definition can be straightforwardly extended to dissipative dynamics as well.
For the problem of extracting quantities related to molecular behavior, many valuable predictions may be made even with errors in the estimated values of $O(t)$, as long as general trends are captured accurately. For example, electron transfer rates are estimated by the decay of donor-state population over time, which can be robust to significant errors in the state populations when the errors are unbiased and transient. Also, approximate classical-digital methods (such as semiclassical approximations) may provide acceptable estimates even if some details of the dynamics are not captured.
{mdframed}
@@
A typical trapped-ion system consists of a linear chain of ions confined in free space. Radio frequency (rf) electromagnetic fields are supplied by the electrodes of the trap, creating a pseudo-potential that constrains the atomic positions. After the ions are sufficiently cooled using lasers, the ions’ interactions with the confinement potential and the Coulomb repulsion among the ions reach equilibrium, producing a linear chain of ions.
A typical ion has many accessible electronic states, and two stable states can be exploited to establish an atom-based qubit with excellent coherence properties. Initialization of the qubit to a desired state, and measurement of the qubit state, may be performed through optical excitation with lasers [49].
Nuclear (bosonic) degrees of freedom exist in the ion chain through its collective vibrating motion. The ion chain’s motion can be described in the language of normal modes, where each mode is approximated by a quantum harmonic oscillator. An $n$-ion chain has $n$ motional modes in each of the ion chain’s axial direction and two other perpendicular (radial) directions, total $3n$ modes. A subset of these modes can interact with the qubits to mediate the entanglement between qubits or to provide the bosonic ingredient for analog simulations. Since the normal modes are spread through the ion chain, each mode can interact with all qubits in the chain (except for qubits located at nodes of the normal mode), allowing flexibility in the structure of the Hamiltonians that can be simulated.
Panel (c) of the figure below visualizes the $n=4$ radial modes of a four-ion chain. The modes are shown in the descending order of mode frequency from top to bottom, and the relative height (direction) of each arrow represents the magnitude (sign) of the eigenvector component of each normal mode. We denote the mode with equal eigenvector components of all ions [top in (c)] as the center-of-mass (CM) mode and all other modes as non-CM modes.
To manipulate the qubit states and motional states of the ion chain, electromagnetic fields are applied using lasers or magnetic-field gradients [49]. Panels (a) and (b) of the figure below show the example of using lasers to couple the leftmost ion’s qubit states to one of the motional modes [third from top in (c)]. The two-photon Raman transition between the qubit states via the excited atomic state $|e\rangle$ is induced using the global (orange) and individually addressed (red and blue) laser beams. The frequency, phase, and amplitude of the laser fields impinging on each ion can be controlled individually, allowing flexibility in defining the Hamiltonian parameters that will be simulated on the ion chain.
The architecture of trapped-ion systems considered in this Review is a single chain of many ions, which scales up to 20-30 ions for state-of-art systems [50, 51]. Another well-known architecture is the quantum charge-coupled device (QCCD) architecture [52], which utilizes operations such as splitting an ion from a chain, shuttling an ion, and merging an ion to another chain. While the QCCD architecture has advantages such as high-fidelity qubit operations, the motional states are not preserved during the splitting and merging, and thus cannot provide the bosonic ingredient for analog simulations.
@@
Assuming that the electronic and vibrational energy states can be obtained, numerically exact methods (see Box 1 for a definition) exist for classical-digital computers to compute quantum molecular dynamics in some regimes. Examples include the multi-configuration time-dependent Hartree (MCTDH) method [56, 57, 58], the Ab Initio Multiple Spawning (AIMS) method [59, 60, 61], the hierarchical equations of motion (HEOM) approach [62, 63, 64], the quasi-adiabatic propagator path integral (QUAPI) method [65, 66, 67, 68, 69, 70, 71], the time-evolving matrix products operators (TEMPO) method [72, 73, 74], the dissipation-assisted matrix product factorization (DAMPF) method [75, 76], and the time-dependent density-matrix renormalization group (tDMRG) method [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. The tDMRG method, for example, can efficiently simulate a composite system of electronic states and bath modes when the composite system may be mapped to a simple configuration, such as one-dimensional chain [79, 80] or star geometry [85] (see Fig. 1d), through an appropriate basis transformation. For the case of one-dimensional chain, large systems (e.g., $N\sim 10^{2}$) with moderate interaction strengths (e.g., $|\kappa_{i,j,k}|\lesssim|\Delta_{i,j}|,\>|E_{i}-E_{j}|,\>\nu_{k}$) can be simulated [80]. However, the tDMRG method is either inaccurate or computationally prohibitive when a transformation to a simple configuration cannot be found, or when the interactions are strong (e.g., $|\kappa_{i,j,k}|\sim$ 10 to $10^{2}$ times $|\Delta_{i,j}|,\>|E_{i}-E_{j}|,\>\nu_{k}$), which is a regime of interest [89, 90, 91], for example, in the study of electron transfer dynamics accompanied by significant nuclear reorganization [92, 24, 93, 91]. Other numerically exact methods also become computationally prohibitive, in terms of computational time or memory demands, when the entanglement among the components becomes significant [94, 73, 67].
@@
This Trotterization-based analog approach to quantum simulation, also suggested for simulating processes in quantum field theory [139], contrasts with the traditional approach in analog trapped-ion simulations, where interactions for each and every term in the simulated Hamiltonian is simultaneously switched on [140, 36]. While the traditional approach enabled various quantum-simulation experiments, even with hardware that does not fully support individual control of qubits (see Ref. 141 for an example of simulating quantum molecular dynamics), the parameters of the simulated Hamiltonian should meet specific constraints [^1] For example, when a quantum spin system is simulated using the traditional approach where all ions are simultaneously illuminated by a monochromatic beam, the coupling strength $\Delta_{i,j}$ between spins $i$ and $j$, mapped to ions $i$ and $j$ indexed from the left to right of the ion chain, should satisfy $\Delta_{i,j}\propto|i-j|^{-\alpha}$ ($30<\alpha<$ 3[140, 141, 36]. that do not exist for the Trotterization-based analog approach.
@@
Using single-qubit rotation (2), spin-dependent force (3), and MS interaction (4), each term in (1) can be simulated. Specifically, the terms $E_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}$ are mapped to (2), the terms $\kappa_{i,i,k}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ are mapped to (3), and the terms $\Delta_{i,j}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{j}$ ($i\neq j$) are mapped to (4). The terms $\kappa_{i,j,k}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{j}(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ ($i\neq j$) are also mapped to a sequence of operations (2)-(4), as described in Ref. 151. The evolution with respect to each term in (1) by the Trotterization time step $\tau/S$ is simulated by choosing the field parameters and time duration of each mapped trapped-ion operation appropriately. By repeating the evolution for all terms and Trotterization steps, the evolution of the entire system described by the LVCM Hamiltonian is simulated.
@@
tDMRG simulations are performed by putting (5) into a one-dimensional topology, where the electronic sites and the $N$ bath modes form a linear chain, using the unitary transformation described in Refs. 80, 85. This strategy can be used to describe the dynamics efficiently for a donor-acceptor system coupled to a bath with hundreds of modes that have a moderately large reorganization energy ($\lambda\lesssim 10\Delta$). However, the computational cost grows as $\kappa\propto\sqrt{\lambda}$ increases, and as the entanglement among the components grows. In contrast, the Ehrenfest simulations can be performed at significantly lower computation cost than the tDMRG simulations, at the expense of accuracy in describing the dynamics only up to a short time compared to $\hslash/\Delta$, as the semiclassical mean-field approximations often break down at longer times.
@@
Figure 1: Comparison of classical (tDMRG and Ehrenfest) and quantum (trapped ion) methods for simulating the model Hamiltonian (5). $\Delta=0.08679$ eV, $\nu_{k}=[0.08679+0.01240(k-1)/(N-1)$] eV for $k=1,..N$, where $N$ is the number of bath modes. For the trapped-ion method, we assume using $(\lceil N/2\rceil+1)$-ion chain to generate $N$ radial non-CM motional modes. (a) Evolution of the donor population, for the indicated reorganization energy values ($\lambda$) and $N=2$, simulated using various methods. For the curves labeled “Trapped ion”, we numerically simulate the evolution of a trapped-ion simulator system. The ideal case assumes no decoherence, and the noisy case assumes state-of-art noise parameters in Table 1. For both cases, shot noise (population uncertainty due to finite number of binary measurements) is not shown. In all panels, the blue and red solid curves overlap. (b) Computational run time for the tDMRG method, with various values of $\lambda/\Delta$ and $N$ [labels shared with (c)]. The additive error for each population value is less than $10^{-3}$ ($10^{-2}$) for the solid (dashed) curves. See Ref. 85 for details of the method (the interaction-picture chain-geometry algorithm). The run time scales nearly exponentially with $\sqrt{\lambda/\Delta}$, where the rate of increase is smaller for the dashed curves. Computational run time for the Ehrenfest method is negligible compared to the tDMRG method. (c) Estimated experimental time for the trapped-ion simulation, with various values of $\lambda/\Delta$ and $N$. We assume 40 equally-spaced time steps and 100 (2500) simulation runs per time step for the left (right) vertical axis, where each run produces a binary measurement result. Dotted line shows the total time spent for cooling, state preparation, and measurement, which is estimated to be 4.25 ms [160] times the total number of simulation runs. The estimated experimental time roughly scales linearly with $\sqrt{\lambda/\Delta}$. (d) A schematic representation of the star-geometry model simulated here, where $N=5$ [see (5)]. The grey circle, denoted ‘S’, represents the system with two electronic energy levels, and each parabola represents a bath mode.
Even when a pulse is resonant with the sideband frequency of a mode ($\mu=\omega_{k}$), unwanted off-resonant excitations of other modes, also known as cross-mode coupling, may occur because of the finite spacings between the mode frequencies. Naïvely, this leads to a constraint that $\tilde{\Omega}_{p,k}$ should be much smaller than the mode-frequency spacing, which imposes a lower bound on the time duration of the corresponding operations. However, cross-mode coupling can be actively removed by modulating the field parameters during the operation. Here, we consider using a frequency($\mu$)-modulated pulse [161, 162] for each spin-dependent-force operation, such that cross-mode coupling is removed using operations with as short a time duration as possible. The required time duration increases as the number of ions increases and mode-frequency spacing decreases. For example, for the case of $\lambda=30\Delta$ and $S=600$, the length of each pulse that simulates a single Trotterization step averaged over all modes is 15.7 $\mu$s, 17.4 $\mu$s, and 19.0 $\mu$s for a 2-ion ($N=2$), 3-ion ($N=3,4$), and 4-ion ($N=5$) chain, respectively, where the number of ions is set as $\lceil N/2\rceil+1$. The relevant system parameters for a state-of-art trapped-ion setup [163] are shown in Table 1.
@@
Table 1: System and noise parameters for the state-of-art trapped-ion system in Refs. 163, 164 that are used in Fig. 1. Motional modes in two radial directions of ion chain consisting of two to four ions are used. Motional-mode coherence time is taken from Ref. 163 and motional-mode heating rate and laser coherence time are taken from Ref. 164. The heating rate of CM motional modes is of the order of 100 quanta/s, so we assume that CM modes are not used in our trapped-ion simulations.
In state-of-art trapped-ion systems, the leading sources of errors are motional-mode dephasing, motional-mode heating, and laser dephasing [163, 164, 165]. To simulate the evolution of a trapped-ion system in the presence of noise, we use Qutip [166] to solve the Lindblad master equation [48], following the method in Ref. 163 which provides reliable predictions of the system performance [163, 164]. We use the best recorded noise parameters from a state-of-art trapped-ion system [163, 164] as shown in Table 1, which are far from the fundamental limits and can, in principle, be improved.
@@
Color gradients in the figure illustrate the qualitative performance (accuracy and computational cost) of three classes of quantum dynamics simulations of LVCMs: numerically exact classical-digital methods, approximate classical-digital methods, and analog trapped-ion simulations on a noisy device. The qualitative performance of these methods is illustrated over several regimes of coupling strength between the electronic states and bath modes, quantified by the ratio of the reorganization energy $\lambda$ to the coupling strength between electronic states $\Delta$.
If $\lambda\ll\Delta$, the numerically exact methods are generally efficient and accurate, since the entanglement among the electronic states and bath modes is weak. In this regime, the approximate methods are also accurate, at least for a few dynamical cycles of the system. This is because the influence of the bath on the electronic states is weak, and errors associated with neglecting non-classical trajectories of the bath modes have little influence on the evolution of the electronic states.
In the regime where $\lambda\gg\Delta$, the numerically exact methods converge slowly and generally perform poorly because of strong entanglement among the electronic states and bath modes. However, the approximate methods generally perform well, providing reliable predictions for the donor-state population decay rate, for example. This acceptable performance arises when the coupling is strong and the bath modes are highly-excited harmonic oscillators that often behave almost classically.
In the intermediate regime where $\lambda\gtrsim\Delta$, the coupling can be sufficiently strong that large entanglement renders the numerically exact methods computationally intractable, yet weak enough that the bath modes are not sufficiently highly excited to behave classically. As such, this regime appears to be a compelling target for testing whether an analog trapped-ion simulation may outperform classical-digital algorithms for simulating quantum molecular dynamics economically, at least in the space of LVCMs. The upper bound of $\lambda/\Delta$ that may allow a trapped-ion simulation to produce sufficiently accurate predictions is determined by the magnitude of the noise relative to the strength of native operations of the trapped-ion system.